# Singular Excitement Beyond the Horizon of a Rotating Black Hole

### v1

@@ III.5 Detector transition rate at the singularity
Figure 6: The transition rate at the singularity (˙$\dot{\mathcal{F}}_{\pi\ell/2}$) as a function of black hole angular momentum, for different black hole masses. Plots (a) and (b) correspond to a detector energy gap of $E\ell=-5$, while plots (c) and (d) correspond to a gap of $E\ell=+5$. For all plots, $\zeta=0$ and $r_{0}/r_{+}=100$. When the black hole mass is large, the transition rate at the singularity is roughly constant for all angular momenta. When black hole mass is small, the transition rate at the singularity decays as an inverse power of $J/M\ell$, as shown by parallel linear traces on the double-log plot (a), (c). However, at intermediate masses, the relationship between ˙$\dot{\mathcal{F}}_{\pi\ell/2}$ and $J/M\ell$ is highly non-trivial, with fluctuations and glitches. Plots (b) and (d) highlight the non-trivial glitch structure within this intermediary band of masses.
@@
As observed in previous sections, the transition rate of the detector does not diverge at the singularity when the black hole is rotating. This absence of a divergence is an important qualitative difference between the static and rotating BTZ scenarios. We conjecture that the divergence in the transition rate for vanishing angular momentum at the time the detector reaches the singularity is a consequence of the manifold breaking down at the singularity. We find for the rotating black hole that the transition rate is everywhere finite, commensurate with the manifold being well defined at the singularity.
@@
In Fig. 6, we show the transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of the detector at the singularity (i.e., when $\tau/\ell=\pi/2$) as a function of black hole angular momentum $J/M\ell\in(0,1)$ for black hole mass ranging from $M=10^{-6}$ to $M=10$. We compare the $E\ell=-5$ and $E\ell=+5$ cases.
@@
When black hole mass is large, the detector’s transition rate at the singularity is essentially independent of the angular momentum $J$. This occurs because the image sum converges more quickly for larger $M$. When $M$ is sufficiently large, the sum is dominated by the $n=0$ term, and thus, in the large mass limit, the transition rate at the singularity tends to the transition rate in pure AdS3 spacetime.
@@
When black hole mass is small, we observe the detector’s transition rate at the singularity to decay as an inverse power of $J/M\ell$, as seen from the straight lines in the double-log plot in Fig. 6 (a), (c). As long as the mass is non-zero, the transition rate at the singularity will be finite. Numerically we find that the transition rate continues to increase as $M\rightarrow 0$ for all $J$.
@@
Between the small and large mass extremes is a range of masses for which the relationship between the transition rate at the singularity and the angular momentum is highly non-trivial. For these masses the transition rate at the singularity is a non-monotonic function of $J/M\ell$. Moreover, there are glitches in the transition rate, which occur when the geodesic distance under one of the square roots in (16) vanishes at $\tau/\ell=\pi/2$. In Fig. 6, we considered the transparent boundary condition $\zeta=0$, and thus the glitches in the plots correspond only to $\sigma_{n}=0$. However, for the other boundary conditions, one obtains additional glitches when $\sigma_{n}+2=0$.
@@
The behaviour of the transition rate as a function of $J/M\ell$ does not depend qualitatively on the gap of the detector. We obtain the same large, small, and intermediary response structure at the singularity for positive detector energy gaps ($E\ell=+5$) and negative ones ($E\ell=-5$). This occurs because the Wightman function depends on $M$ but not on $E$, whereas $E$ controls the phase in (16). Thus, the small and large mass limiting behaviours are independent of $E$.
@@ IV Conclusion
We numerically calculated the transition rate of an Unruh-DeWitt detector coupled to a massless conformal scalar field as the detector falls freely into a rotating BTZ black hole along a co-rotating trajectory. We explored the effects of different black hole masses, black hole angular momenta, detector energy gaps, and field boundary conditions at infinity. To calculate the transition rate, we expressed the Wightman function of the BTZ black hole as an image sum of Wightman functions in AdS3, and then computed the transition rate as a sum of the contributions arising from each image. Unlike in the non-rotating case, the terms in the transition rate when rotation is present are not invariant under $n\rightarrow-n$. We found that selecting different boundary condition for the field at infinity did not change the qualitative behaviour of the transition rate curves. However, black hole mass and detector energy gap did significantly impact the transition rate, while black hole angular momentum also played a role. Specifically, the presence of angular momentum caused a stretching of the transition rate curve away from the singularity.
@@
As with the non-rotating BTZ case, we discovered that the detector’s transition rate is non-differentiable at certain discrete values of the detector’s proper time. Following the nomenclature in [14], we refer to these points of non-differentiability as glitches. There are four types of glitches in the rotating BTZ case, which are asymmetric for $n\rightarrow-n$. The exotic glitches (i.e., those that were not seen in the static case) live entirely behind the inner horizon, and thus vanish as expected as angular momentum of the black hole goes to zero. The locations of the glitches depend on the black hole mass and angular momentum, and also (though we did not consider it in this study) the initial radial position of the detector.
@@
We observed that, in the presence of rotation, the detector’s transition rate at the singularity does not diverge. Furthermore, we identified a range of masses for which the transition rate at the singularity exhibits glitches when plotted as a function of black hole angular momentum. In the small mass limit, the transition rate at singularity decays as an inverse power of angular momentum, while in the large mass limit, the transition rate at singularity tends to the transition rate in pure AdS3 spacetime.
@@
Our study was motivated firstly by the observation that the response function of a detector interacting with the Hartle-Hawking(-Israel) state of a massless scalar field and freely falling into a (3+1)-dimensional Schwarzschild black hole exhibits a local extremum at the horizon crossing [13]. Following this work with the Schwarzschild black hole was a further study [14], wherein a UDW detector was permitted to fall radially into a (2+1)-dimensional non-rotating BTZ black hole. In this latter study, in addition to the discovery of glitches, it was found that more than one local extremum could exist near the horizon under certain parameters. Our results for the (2+1)-dimensional rotating BTZ black hole are commensurate with both of these previous studies, and moreover provides a more general description of UDW detectors in BTZ spacetime (from which the previous non-rotating BTZ case can be derived). The possibility of using glitches as an early warning for horizon crossing, so long as black hole mass is sufficiently small, remains valid in the presence of rotation.
@@
A future avenue of research would be to study the extanglement dynamics of multiple detectors in rotating BTZ spacetime, specifically when one or more of the detectors falls behind the horizon. Additionally, the work on infalling UDW detectors can be extended to other black hole spacetimes where the curvature may not be constant. The role of hidden topologies may also be a topic of interest–for instance, comparing the transition rates of detector falling into a BTZ black hole versus its geon counterpart. Finally, one could explore the transition probability of a detector falling into a black hole while coupled to other fields–for example, a fermionic field.
@@ Appendix A Individual terms
Figure A1: A few individual contributions to the transition rate calculated with $\zeta=0$ and the same parameters as in Fig. 5. (a), (c) and (e) correspond to terms of negative $n$, meanwhile (b), (d) and (f) correspond to terms of positive $n$. The time at the outer (inner) horizon is indicated with a solid (dashed) vertical line. The glitches are shown with markers following the notation in Fig. 5. The circle and square markers correspond to the first and second type of glitches which arise from the terms independent of the boundary condition.
@@
Figure A2: Individual contributions to the transition rate calculated with $\zeta=-1$ or $\zeta=1$ and the same parameters as in Fig. 5. We show values of $n$ that differ considerably with respect to the terms calculated with $\zeta=0$. These correspond to (c), (e) and (f) of Fig. A1. The glitches are shown with markers following the notation in Fig. 5. The diamond and triangle markers correspond to the third and fourth type of glitches, which arise from singularities in the term that depend on the field’s boundary conditions.
@@
As seen in Fig. A1 (a) and (b), the shape of the first few terms is mostly determined by the oscillations following the first type of glitches. On the other hand, there are no oscillations following the second type of glitches, which are located very close to the time to the singularity. For these first terms, the glitches for negative and positive terms occur at times that are relatively close to each other. As a result, the shape of the individual contributions to the transitions rate are essentially the same for positive and negative $n$ terms.
@@
As observed in Fig. A1 (c) and (d), it is for $n=-5$ and $n=5$ that qualitative differences in the contributions become apparent: while the negative $n$ terms attain three local extrema, the positive $n$ terms attain only two or none. This symmetry loss happens because the glitches of the first type are located in the band between the time to the horizon and the time to the singularity for negative $n$ but are located past the time to the singularity when $n$ is positive. Moreover, these glitches tend to the inner horizon crossing as $n\to-\infty$, which is observed in Fig. A1 (e), but tend to a reflection of the inner horizon crossing about $\pi/2$ for $n\to\infty$.
@@
In Fig. A1 (e) and (f), we observe that the amplitude of the contributions decreases as $|n|$ increases. The terms decaying fast enough is what ensures that the transition rate remains finite over all the domain of the total detection time, including the time to the singularity and the horizons. We also observe that the overall shape is maintained but shrunken along the total detection time axis for bigger $|n|$. This behaviour is what produces the jagged region in the transition rate behind the inner horizon crossing.
@@
The role of the third (diamond) and fourth (triangle) types of glitches is illustrated in Fig. A2, where we have chosen to show the plots analogous to Fig. A1 (c), (e), and (f) because they differ the most when compared to the contributions calculated with $\zeta=0$. For small $|n|$ or for positive $n$, glitches of the third type only occur after the time to the singularity. Hence, these contributions do not differ substantially from $\zeta=0$. However, as the magnitude of $n$ increases, the third type of glitches starts to appear in the contributions for negative $n$ terms, as observed in Fig. A2 (a) and (b). As expected, the oscillation behind this type of glitches has an opposite phase for $\zeta=\pm 1$.
@@
We observe in Fig. A2 (c) and (d) that glitches of both the third and fourth types are present in the terms for negative and large magnitude $n$, but the amplitude of the curves after the fourth type of glitches is negligible compared to the rest of the contribution. Since the glitches of the third type tend to a reflection of the inner horizon crossing with respect to $\pi/2$ as $n\to\infty$, they are not present in the plots of Fig. A2 (e) and (f). Although the behaviour is qualitatively different for $\zeta=\pm 1$, the amplitude of the contributions after the fourth type of glitches is again negligible when compared to other terms.
@@
Overall, we note that the terms that are mostly determined by the first type of glitches, which are the ones with small $|n|$, resemble those that were present in the non-rotating case of [14]. We also note that when more terms are included in the image sum, the difference between the static and the rotating black holes becomes apparent because the exotic types of glitches are present. These new types of glitches, that were absent in the static case, occur close to one another behind the inner horizon crossing and do not pile up at the time to the singularity, which results in transition rate curves that are jagged behind the inner horizon crossing and finite at the time to the singularity.

### v2

@@ III.5 Detector transition rate approaching the singularity
Figure 6: The transition rate ˙$\dot{\mathcal{F}}_{\pi\ell/2}$ as a function of black hole angular momentum, for different black hole masses. Plots (a) and (b) correspond to a detector energy gap of $E\ell=-5$, while plots (c) and (d) correspond to a gap of $E\ell=+5$. For all plots, $\zeta=0$ and $r_{0}/r_{+}=100$. When the black hole mass is large, the transition rate at $\Delta\tau/\ell=\pi/2$ is roughly constant for all angular momenta. When the black hole mass is small, the transition rate at $\Delta\tau/\ell=\pi/2$ decays as an inverse power of $J/M\ell$, as shown by parallel linear traces on the double-log plot (a), (c). However, at intermediate masses, the relationship between ˙$\dot{\mathcal{F}}_{\pi\ell/2}$ and $J/M\ell$ is highly nontrivial, with fluctuations and glitches. Plots (b) and (d) highlight the nontrivial glitch structure within this intermediary band of masses.
@@
As observed in previous sections, the transition rate of the detector does not diverge as it approaches the singularity when the black hole is rotating. This absence of a divergence is an important qualitative difference between the static and rotating BTZ scenarios, and arises because the detector’s point of closest approach is $r=\frac{r_{-}r_{+}}{r_{0}}$ at $\Delta\tau/\ell=\pi/2$.
@@
We conjecture that the divergence in the transition rate for vanishing angular momentum at the time the detector reaches the singularity is a consequence of the manifold breaking down at the singularity. We find for the rotating black hole that the transition rate is everywhere finite, commensurate with the manifold being well defined at the singularity. In Fig. 6, we show the transition rate ˙$\dot{\mathcal{F}}_{\tau}$ of the detector when $\Delta\tau/\ell=\pi/2$ as a function of black hole angular momentum $J/M\ell\in(0,1)$ for black hole mass ranging from $M=10^{-6}$ to $M=10$. We compare the $E\ell=-5$ and $E\ell=+5$ cases.
@@
When the black hole mass is large, the detector’s transition rate at $\Delta\tau/\ell=\pi/2$ is essentially independent of the angular momentum $J$. This occurs because the image sum converges more quickly for larger $M$. When $M$ is sufficiently large, the sum is dominated by the $n=0$ term, and thus, in the large mass limit, the transition rate at $\Delta\tau/\ell=\pi/2$ tends to the transition rate in pure AdS3 spacetime.
@@
When the black hole mass is small, the detector’s transition rate at $\Delta\tau/\ell=\pi/2$ decays as an inverse power of $J/M\ell$, as seen from the straight lines in the double-log plot in Fig. 6 (a), (c). As long as the mass is nonzero, the transition rate remains finite. Numerically we find that the transition rate continues to increase as $M\rightarrow 0$ for all $J$.
@@
Between the small and large mass extremes is a range of masses for which the relationship between the transition rate at $\Delta\tau/\ell=\pi/2$ and the angular momentum is highly nontrivial. For these masses the transition rate ˙$\dot{\mathcal{F}}_{\pi\ell/2}$ is a nonmonotonic function of $J/M\ell$. Moreover, there are glitches in the transition rate, which occur when the geodesic distance under one of the square roots in (16) vanishes at $\tau/\ell=\pi/2$. In Fig. 6, we considered the transparent boundary condition $\zeta=0$, and thus the glitches in the plots correspond only to $\sigma_{n}=0$. However, for the other boundary conditions, one obtains additional glitches when $\sigma_{n}+2=0$.
@@
The behavior of the transition rate as a function of $J/M\ell$ does not depend qualitatively on the gap of the detector. We obtain the same large, small, and intermediary response structure at closest approach to the singularity for positive detector energy gaps ($E\ell=+5$) and negative ones ($E\ell=-5$). This occurs because the Wightman function depends on $M$ but not on $E$, whereas $E$ controls the phase in (16). Thus, the small and large mass limiting behaviors are independent of $E$.
Understanding what happens to the rate at $r=0$ in the rotating case will require placing the detector in non-geodesic motion, a subject for future study.
@@ IV Conclusion
We numerically calculated the transition rate of an Unruh-DeWitt detector coupled to a massless conformal scalar field as the detector falls freely into a rotating BTZ black hole along a corotating trajectory. We explored the effects of different black hole masses, black hole angular momenta, detector energy gaps, and field boundary conditions at infinity. To calculate the transition rate, we expressed the Wightman function of the BTZ black hole as an image sum of Wightman functions in AdS3, and then computed the transition rate as a sum of the contributions arising from each image. Unlike in the nonrotating case, the terms in the transition rate when rotation is present are not invariant under $n\rightarrow-n$. We found that selecting a different boundary condition for the field at infinity did not change the qualitative behavior of the transition rate curves. However, black hole mass and detector energy gap did significantly impact the transition rate, while black hole angular momentum also played a role. Specifically, the presence of angular momentum caused a stretching of the transition rate curve away from the singularity.
@@
As with the nonrotating BTZ case, we discovered that the detector’s transition rate is nondifferentiable at certain discrete values of the detector’s proper time. Following the nomenclature in [14], we refer to these points of nondifferentiability as glitches. There are four types of glitches in the rotating BTZ case, which are asymmetric for $n\rightarrow-n$. The exotic glitches (i.e., those that were not seen in the static case) live entirely behind the inner horizon and thus vanish as expected as the angular momentum of the black hole goes to zero. The locations of the glitches depend on the black hole mass and angular momentum, and also (though we did not consider it in this study) the initial radial position of the detector.
@@
We observed that, in the presence of rotation, the detector’s transition rate at $\Delta\tau/\ell=\pi/2$, at which the detector reaches $r=\frac{r_{-}r_{+}}{r_{0}}$ (its point of closest approach to the singularity) does not diverge. Furthermore, we identified a range of masses for which the transition rate at this point exhibits glitches when plotted as a function of black hole angular momentum. In the small mass limit, the transition rate at $\Delta\tau/\ell=\pi/2$ decays as an inverse power of angular momentum, while in the large mass limit, the transition rate tends to the transition rate in pure AdS3 spacetime.
@@
Our study was motivated first by the observation that the response function of a detector interacting with the Hartle-Hawking(-Israel) state of a massless scalar field and freely falling into a (3+1)-dimensional Schwarzschild black hole exhibits a local extremum at the horizon crossing [13]. Following this work with the Schwarzschild black hole was a further study [14], wherein a UDW detector was permitted to fall radially into a (2+1)-dimensional nonrotating BTZ black hole. In this latter study, in addition to the discovery of glitches, it was found that more than one local extremum could exist near the horizon under certain parameters. Our results for the (2+1)-dimensional rotating BTZ black hole are commensurate with both of these previous studies and, moreover, provide a more general description of UDW detectors in BTZ spacetime (from which the previous nonrotating BTZ case can be derived). The possibility of using glitches as an early warning for horizon crossing, so long as black hole mass is sufficiently small, remains valid in the presence of rotation.
@@
A future avenue of research would be to study the entanglement dynamics of multiple detectors in rotating BTZ spacetime, specifically when one or more of the detectors falls behind the horizon. Additionally, the work on infalling UDW detectors can be extended to other black hole spacetimes where the curvature may not be constant. The role of hidden topologies may also be a topic of interest—for instance, comparing the transition rates of a detector falling into a BTZ black hole versus its geon counterpart. Finally, one could explore the transition probability of a detector falling into a black hole while coupled to other fields—for example, a fermionic field.
@@ Appendix A Individual terms
Figure A1: A few individual contributions to the transition rate calculated with $\zeta=0$ and the same parameters as in Fig. 5. (a), (c), and (e) correspond to terms of negative $n$; meanwhile (b), (d), and (f) correspond to terms of positive $n$. The time at the outer (inner) horizon is indicated with a solid (dashed) vertical line. The glitches are shown with markers following the notation in Fig. 5. The circle and square markers correspond to the first and second types of glitches, which arise from the terms independent of the boundary condition.
@@
Figure A2: Individual contributions to the transition rate calculated with $\zeta=-1$ or $\zeta=1$ and the same parameters as in Fig. 5. We show values of $n$ (corresponding to (c), (e) and (f) of Fig. A1) that differ considerably with respect to the terms calculated with $\zeta=0$. Plots (a), (c), and (e) in this figure depict the $\zeta=-1$ case while (b), (d), and (f) depict the $\zeta=1$ case. The glitches are shown with markers following the notation in Fig. 5. The diamond and triangle markers correspond to the third and fourth types of glitches, which arise from singularities in the term that depend on the field’s boundary conditions.
@@
As seen in Fig. A1 (a) and (b), the shape of the first few terms is mostly determined by the oscillations following the first type of glitches. On the other hand, there are no oscillations following the second type of glitches, which are located very close to $\Delta\tau/\ell=\pi/2$. For these first terms, the glitches for negative and positive terms occur at times that are relatively close to each other. As a result, the shape of the individual contributions to the transition rate is essentially the same for positive and negative $n$ terms.
@@
As observed in Fig. A1 (c) and (d), it is for $n=-5$ and $n=5$ that qualitative differences in the contributions become apparent: while the negative $n$ terms attain three local extrema, the positive $n$ terms attain only two or none. This symmetry loss happens because the glitches of the first type are located in the band between the time to the horizon and the time to $r=\frac{r_{-}r_{+}}{r_{0}}$ for negative $n$ but are located past this time when $n$ is positive. Moreover, these glitches tend to the inner horizon crossing as $n\to-\infty$, which is observed in Fig. A1 (e), but tend to a reflection of the inner horizon crossing about $\pi/2$ for $n\to\infty$.
@@
In Fig. A1 (e) and (f), we observe that the amplitude of the contributions decreases as $|n|$ increases. The terms decay quickly enough to ensure that the transition rate remains finite over the entire domain of the total detection time, including the time at closest approach to the singularity and the horizons. We also observe that the overall shape is maintained but shrunken along the total detection time axis for bigger $|n|$. This behavior produces the jagged region in the transition rate behind the inner horizon crossing.
@@
The role of the third (diamond) and fourth (triangle) types of glitches is illustrated in Fig. A2, where we have chosen to show the plots analogous to Fig. A1 (c), (e), and (f) because they differ the most when compared to the contributions calculated with $\zeta=0$. For small $|n|$ or for positive $n$, glitches of the third type only occur after $\Delta\tau/\ell=\pi/2$. Hence, these contributions do not differ substantially from $\zeta=0$. However, as the magnitude of $n$ increases, the third type of glitches starts to appear in the contributions for negative $n$ terms, as observed in Fig. A2 (a) and (b). As expected, the oscillation behind this type of glitches has an opposite phase for $\zeta=\pm 1$.
@@
We observe in Fig. A2 (c) and (d) that glitches of both the third and fourth types are present in the terms for negative and large magnitude $n$, but the amplitude of the curves after the fourth type of glitches is negligible compared to the rest of the contribution. Since the glitches of the third type tend to a reflection of the inner horizon crossing with respect to $\pi/2$ as $n\to\infty$, they are not present in the plots of Fig. A2 (e) and (f). Although the behavior is qualitatively different for $\zeta=\pm 1$, the amplitude of the contributions after the fourth type of glitch is again negligible when compared to other terms.
@@
Overall, we note that the terms that are mostly determined by the first type of glitch, which are the ones with small $|n|$, resemble those that were present in the nonrotating case of [14]. We also note that when more terms are included in the image sum, the difference between the static and the rotating black holes becomes apparent because the exotic types of glitches are present. These new types of glitches, which were absent in the static case, occur close to one another behind the inner horizon crossing and do not pile up at $\Delta\tau/\ell=\pi/2$, which results in transition rate curves that are jagged behind the inner horizon crossing and finite at $\Delta\tau/\ell=\pi/2$.